Carol puts some green and red unit cubes in a box. The ratio of the number of green cubes to the number of red cubes is 2:1. She adds 12 more red cubes in the box and the ratio becomes 4:5.
a) How many green cubes are there in the box?
b) How many red cubes does Carol have in the end?

Answer by scott8148(6628)

"The ratio of the number of green cubes to the number of red cubes is 2:1"
g = 2r

"She adds 12 more red cubes in the box and the ratio becomes 4:5"
5g = 4(r + 12) ___ 5g = 4r + 48

substituting ___ 5(2r) = 4r + 48 ___ 10r = 4r + 48 ___ 6r = 48 ___ r = 8

(a) g = 2(8) ___ g = 16

(b) r + 12 = 8 + 12 = 20